OpenStudy (anonymous):

Write the equation of the line that is parallel to the line y = 2x + 2 and passes through the point (5, 3). y = 2x - 7 y = - one halfx - 2 y = 2x - 2 y = - one halfx - 7

OpenStudy (anonymous):

do u know whats point slope intercept form?

OpenStudy (anonymous):

y - y1 = m(x - x1) ever heard of this ?

OpenStudy (anonymous):

A line parallel to y = 2x+2 has the same slope. Use y-(y1)=m(x-x1), and using y1= 3, and x1 = 5 to solve for y

OpenStudy (anonymous):

okay so first of all if its parallel therefore the slope will remain the same for the other equation too

OpenStudy (anonymous):

so y-3=2(x-5)

OpenStudy (anonymous):

so substitute the given points in the point slope intercept form and the slope(m)=2

OpenStudy (anonymous):

yes u got it

OpenStudy (anonymous):

now how do i solve

OpenStudy (anonymous):

multiply 2 with (x - 5)

OpenStudy (anonymous):

which is going to be 2x - 10

OpenStudy (anonymous):

+3

OpenStudy (anonymous):

then add 3 to both the sides

OpenStudy (anonymous):

which will give u y=2x - 7

OpenStudy (anonymous):

now i get it thanks a lot

OpenStudy (anonymous):

no prb

OpenStudy (anonymous):

my pleasure

OpenStudy (anonymous):

can you help me with this one Select the coordinates of two points that are on the line x = 5. (5, 0) and (0, 5) (5, 0) and (5, -2) (5, 0) and (0, 2) (0, 5) and (0, 2)

OpenStudy (anonymous):

let me see

OpenStudy (anonymous):

should be the second one as both the coordinates have 5 in the x - place

OpenStudy (anonymous):

so the two last ans cant be it

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

and also the first one can't be the ans.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

and this one Is the line for the equation y = - 8 horizontal or vertical? What is the slope of this line? Select the best answer. The line is horizontal. The slope is undefined. The line is horizontal. The slope is 0. The line is vertical. The slope is undefined. The line is vertical. The slope is 0.